OpenStudy (anonymous):

A(n)=-7.6+(n-1)(3) Please help

OpenStudy (anonymous):

@ganeshie8 @Preetha

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

Find the 4,6 and 13 terms of the sequence

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

im her

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

go their it can help you

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

4th term = A(4) =-7.6+3(4-1) = ? 6th term = A(6) = -7.6+3(6-1) = ? 13th term = A(13) = -7.6+3(13-1) = ?

OpenStudy (anonymous):

4th: A=.35? @tom982

OpenStudy (anonymous):

\[A(4)=-7.6+3(4-1)=-7.6+3(3)=-7.6+9=1.4\]